1. Intrigration of dx/1+e^2x
1
Expert's answer
2021-10-14T04:27:44-0400
"\int\dfrac{1}{e^{2x}+1}dx=\int\dfrac{e^{-2x}}{1+e^{-2x}}dx"

"u=1+e^{-2x}, du=-2e^{-2x}dx"

"\int\dfrac{1}{e^{2x}+1}dx=\int\dfrac{e^{-2x}}{1+e^{-2x}}dx=-\dfrac{1}{2}\int\dfrac{1}{u}du"

"=-\dfrac{1}{2}\ln(|u|)+C=-\dfrac{1}{2}\ln(1+e^{-2x})+C"

"\int\dfrac{1}{e^{2x}+1}dx=-\dfrac{1}{2}\ln(1+e^{-2x})+C"

"=x-\dfrac{1}{2}\ln(1+e^{2x})+C"